# Generic motives and motivic cohomology of fields

### v1

@@ 6. Categorical presentation of cycle modules
. We let ~$\tilde{\mathscr{E}}_{k}$ be the category whose objects are given by couples $(E,n)$ where $E/k$ is a function field, and $n\in\mathbb{Z}$ an integer. The morphisms of ~$\tilde{\mathscr{E}}_{k}$ are abelian groups defined by the following generators $D*$ and relations $R*$ as displayed below, where maps are always assumed to be $k$-algebra morphisms, and the valuations are always understood as (geometric) valuation of function fields:
@@
    Let $\varphi:E\rightarrow L$, $v$ a valuation on $L/k$ wich is zero on $E^{\times}$: $\partial_{v}\circ\varphi_{*}=0$.
@@
    Let $(E,v)$ be a valued function field over $k$ and $\pi$ uniformizer on $v$: $\partial_{v}\circ\gamma_{\{-\pi\}}\circ\varphi_{*}={\bar{\varphi}}_{*}$.
@@
According to [Ros96, Def. 1.1], a cycle premodule over $k$ is simply a (covariant) functor $M:\tilde{\mathscr{E}}\rightarrow\mathscr{A}b$.
6.0.2.
@@
Let $F$ be field with an arbitrary valuation (non necessarily discrete)
@@
Recall that the rank of $v$ is the rank of the abelian group $\mathrm{Im}(v)$ (loc. cit. Definition after Th. 1.7), or equivalently the (Krull) dimension of the valuation ring $\mathcal{O}_{v}$ (loc. cit. Corollary after Th. 1.7). If $v$ has finite rank $r$, then $\Gamma\simeq\mathbb{Z}^{r}$ and $v$ can be defined as a composition of $r$ valuations $(v_{1},\cdots,v_{r})$, $v_{i+1}$ being a valuation on the residue field $\kappa_{v_{i}}$ (see loc. cit. Rem. 1.7).
6.0.3.
@@
Following [Ros96, Rem. 1.6], we define a $ZZ$-graded ring associated to $v$ by the formula:
@@
Considering the obvious projection map $p:\operatorname{K}^{M}_{*}(F)\rightarrow\operatorname{K}^{M}_{*}(v)$, there exists a morphism $i:\operatorname{K}^{M}_{*}(\kappa_{v})\rightarrow\operatorname{K}^{M}_{*}(v)$ of graded ring uniquely defined by the universal property [^43] This follows from the exact sequence (6.0.3.a) by induction on the degree of symbols in Milnor K-theory:
@@
$\bigoplus_{r\geq 0,(v_{*},L)\in\mathcal{I}_{r}(F,E)}\operatorname{K}^{M}_{*}(L)\otimes_{\operatorname{K}^{M}_{*}(\kappa_{v})}\operatorname{K}^{M}_{*}(v)$ $\rightarrow\operatorname{Hom}_{\tilde{\mathscr{E}}_{k}}\big{(}(F,0),(L,*)\big{)}$
@@
Obviously, given that the functor $(E,0)\mapsto(E,1)$ is an auto-equivalence of the category ~$\tilde{\mathscr{E}}_{k}$, this completely describes the morphism in the latter category. The proof follows by applying the relations (R*) to put morphisms in ~$\tilde{\mathscr{E}}_{k}$ in the above normalized form. A key point is Lemma 1.9 of [Ros96].
7. Recall on Bousfield localizations
Recall the following definition due to Bousfield (originally phrased in terms of model category [Bou75]):
Definition 7.0.1.
Let $R$ be a ring. A map $f:X\rightarrow Y$ in the homotopy category $\operatorname{\mathscr{H}}$, will be called an $R$-equivalence if the induced map $f_{*}:H_{*}(X,R)\rightarrow H_{*}(Y,R)$ is an isomorphism. An object $T$ of $\operatorname{\mathscr{H}}$ is called $R$-local if for any $R$-equivalence $f:X\rightarrow Y$, the induced map $f_{*}:Map(Y,P)\rightarrow\operatorname{Map}(X,P)$ is an equivalence.
We let $\operatorname{\mathscr{H}}_{R}$ be the localization of the $\infty$-category $\operatorname{\mathscr{H}}$ with respect to $R$-equivalences, and call it the Bousfield $R$-local homotopy category. This is a presentable $\infty$-category, that can be identified with the full $\infty$-subcategory of $\operatorname{\mathscr{H}}$ spanned by $R$-local objects (see [Lur09]). Moreover, there exists an adjunction of $\infty$-categories
$\pi_{R}:\operatorname{\mathscr{H}}\leftrightarrows\operatorname{\mathscr{H}}_{R}:L^{\prime}_{R}$
It is customary to let $L_{R}=L^{\prime}_{R}\circ\pi_{R}$ and call this the $R$-localization functor. For any space $X$, there exists a canonical map
$X\rightarrow L_{R}(X)$
which is an $R$-equivalence and such that $L_{R}(X)$ is $R$-local. This map is universal.
Example 7.0.2.
  1. (1)
    When $R=\mathbb{Z}$, $L_{\mathbb{Z}}(X)=X^{+}$ for any space $X$ where $X^{+}$ denotes Quillen’s plus construction. This actually follows from the construction of Quillen: see [Hoy].
  2. (2)
    When $R=\mathbb{Q}$, and $X$ is simply connected $L_{\mathbb{Q}}(X)$ can be computed as either the Sullivan or the Quillen model of rationalization of simply connected spaces. See [Iva22].

### v3

@@ 7. Categorical presentation of cycle modules
[…]
@@
    Let $\varphi:E\rightarrow L$, $v$ a valuation on $L/k$ which is zero on $E^{\times}$: $\partial_{v}\circ\varphi_{*}=0$.
@@
    Let $(E,v)$ be a valued function field over $k$ and $\pi$ uniformizer of $v$: $\partial_{v}\circ\gamma_{\{-\pi\}}\circ\varphi_{*}={\bar{\varphi}}_{*}$.
@@
According to [Ros96, Def. 1.1], a cycle premodule over $k$ is simply a (covariant) functor $M:\tilde{\mathcal{E}}\rightarrow\mathrm{A}b$.
7.0.2.
@@
Let $F$ be field with an arbitrary valuation (not necessarily discrete)
@@
Recall that the rank of $v$ is the rank of the abelian group $\mathrm{Im}(v)$ (loc. cit. Definition after Th. 1.7), or equivalently the (Krull) dimension of the valuation ring $\mathcal{O}_{v}$ (loc. cit. Corollary after Th. 1.7). If $v$ has finite rank $r$, then $\Gamma\simeq\mathbb{Z}^{r}$ and $v$ can be defined as a composition of $r$ valuations $(v_{1},\cdots,v_{r})$, $v_{i+1}$ being a valuation on the residue field $\kappa_{v_{i}}$ (see loc. cit. Remark 1.7).
7.0.3.
@@
Following [Ros96, Rem. 1.6], we define a $\mathbb{Z}$-graded ring associated to $v$ by the formula:
@@
Considering the obvious projection map $p:\operatorname{K}^{M}_{*}(F)\rightarrow\operatorname{K}^{M}_{*}(v)$, there exists a morphism $i:\operatorname{K}^{M}_{*}(\kappa_{v})\rightarrow\operatorname{K}^{M}_{*}(v)$ of graded rings uniquely defined by the universal property [^75] This follows from the exact sequence (7.0.3.a) by induction on the degree of symbols in Milnor K-theory:
@@
$\bigoplus_{r\geq 0,(v_{*},L)\in\mathcal{I}_{r}(F,E)}\operatorname{K}^{M}_{*}(L)\otimes_{\operatorname{K}^{M}_{*}(\kappa_{v})}\operatorname{K}^{M}_{*}(v)$ $\rightarrow\operatorname{Hom}_{\tilde{\mathcal{E}}_{k}}\big{(}(F,0),(L,*)\big{)}$
@@
[…]